OpenStudy (moongazer):

What percent profit was made on a pair of pants that was bought for $600 and sold for $630? solution plsss^_^

OpenStudy (anonymous):

im working on it

OpenStudy (anonymous):

600/30=20, so you make 20% profit

OpenStudy (moongazer):

the answer is 5% but i don't know how it become 5

OpenStudy (anonymous):

You make a 5% profit based on the amount of money you actually made. Profit = Revenue - Cost Profit = 630 - 600 Profit = $30. Now you are looking for the percentage of the profit so you what percent of 600 is 30. $30/600 = .05 which equals 5%